# Electromagnetic energy calibration of the SoLid detector with horizontal muons

### v3

@@ 4 Relative calibration
The System Matrix is first filled with the initial values of the projectors according to Equation 2.1. For each fibre, the projection is computed by multiplying the number of photons arriving () by the MPPC gain (). The gain for the entire SoLid detector was equalised at the beginning of each data-taking period [11]. Therefore, the projector is defined as:
@@ 4.1 KL divergence
The Kullback–Leibler (KL) divergence method [12] is a type of statistical distance (or dissimilarity test), which quantifies how much a probability distribution is different from a reference probability distribution . In the case of the SoLid experiment, both distributions must obey the Poisson law; hence, the KL divergence, slightly modified with respect to [12], is given by:
This class of equations is solved with a Lambert function [13]. This is a multivalued function with different branches, which works only for certain values of . In our implementation, is either 22.5% (main cube with four fibres) or 1.25% (adjacent cubes, 8 fibres). The parameter is the mean of a Gaussian model fit to the initial distribution of . Equation 4.2 is solved for (mean value of the measured distribution) as follows:
As a result, it is not the distribution of values that is fit. Instead, we fit the statistical distance for each set of projector values from the Poisson law. The outliers in the transformed distribution are much more pronounced and can be straightforwardly excluded from the fit. The result of the application of the KL divergence method is shown in Figure 8. As such, the total sum of the light-sharing fractions from the four fibres of the main cube and eight fibres of the adjacent cubes improves from 96% to 99%.
@@ 4.2 Homogenisation of the response
The second stage of the relative calibration is the homogenisation of the response from individual cubes. The characteristics of the detection units differ, causing a variation in the response of the detector. The energy loss of the muon serves as a calibration reference. The sum of the responses of the 12 fibres involved within the plane is considered as the total energy deposit made by the muon while crossing it (as in the variable from Equation 4.1). In addition, all muon candidates are selected from Type 2 muons, for which there is a successfully reconstructed track. The path length of the muon within a certain plane is determined from this track information. Thus, the energy loss is estimated as:
The energy loss distribution for a given cube is obtained by applying Equation 4.4 to the subset of muons that cross it. A muon is rejected if the cube is not reconstructed with 4 fibres (except for the cases where the fibre is permanently dead). This additional selection compensates for the dying fibre effect presented in Section 3.1. The result of the selection is shown in Figure 9. A Landau function convoluted with a Gaussian is fit to the resulting energy loss distribution. The model has 3 free parameters: the Gaussian width, the Landau scale factor, and its most probable value (, where shows that the fit result comes from the energy deposit ). The latter is used as the denominator in the homogenisation procedure.
The mean detector response is used as the numerator. It is denoted as and is calculated as the mean value of the distribution with all the values of of the individual detection cubes. Projector values are weighted with respect to the average detector response. This re-weighting concludes the relative calibration of the detector and establishes the final values of the elements of the System Matrix. The impact of the homogenisation procedure is, as expected, the most pronounced for planes with permanently dead fibres. As an illustration, Figure 10 displays the response of the 4 fibres of each main cube in plane 48. The dark blue area (half of the plane) corresponds to the dead fibres. Even under such conditions, the relative calibration works reasonably well and allows the events appearing in that region of the detector to be adequately reconstructed.
@@ 5 Absolute calibration
The results of the relative calibration must be complemented by the absolute energy scale. This is a factor that converts the energy measurement from units provided by the electronics (ADC) to physics units (MeV). One possibility of determining it consists of using the energy-loss distribution from the cosmic horizontal muon. Since the relative calibration provides equalised responses from all the individual detection units, it is sufficient to compute a single conversion factor instead of proceeding cube-by-cube. All the energy loss values are therefore gathered in a single distribution, to which the Landau function convoluted with the Gaussian is fit. The most probable value of Landau gives the reference value in ADC units 111technically in pixel avalanches, or PA, but since the gain is equalised for all the fibres, the conversion factor from PA to ADC is known. The value in MeV is obtained from a dedicated Geant4 simulation. It includes the full-scale detector geometry that is combined with the CRY engine [14] to simulate muons. All energy loss values obtained from the simulation are also combined in a single distribution. The ratio between the most probable values extracted from two fits defines the absolute energy scale.
Figure 11 shows the distributions of the energy losses values obtained with muons in three different energy regimes (0.5, 5, and 50 GeV). Variations in the average response are at the level of 10%. This provides the fundamental limit to the precision of this method. Any improvement would require a more accurate knowledge of the energy distribution of horizontal muons.
@@ 5.1 Absolute calibration with the americium-beryllium source
A novel method (to our knowledge) has been introduced by making use of an americium-beryllium (AmBe) radioactive calibration source that was used in SoLid to calibrate the neutron signal. The idea consists of measuring the response of the detector to the 4.44 MeV gamma emitted by excited 12C, with a branching fraction of 60%. These gammas can experience an electron-positron () pair conversion which provides a well-defined monoenergetic calibration source of 3.4 MeV. Although Compton scattering is the dominant process, the conversion appears in 3.5% of the cases according to a generic Geant4 simulation (with the G4EmStandardPhysics physics list). The presence of the positron makes the calibration signature identical to that of the IBD signal candidates. Therefore, all tools developed for the topological identification of IBD events are highly relevant and can be used to select .
The chain that produces the production of 4.4 MeV gammas is as follows:
@@
The initial information on the electromagnetic part of the signal provided by AmBe is obtained from a dedicated Geant4 simulation. It consists of a simulation of the full-scale SoLid detector geometry, where the radioactive source is placed in one of the positions available between the modules. In order to optimise the computational cost of the process, only and are simulated directly, instead of the entire chain shown in 5.1. The simulation indicates that the pair deposits its energy predominantly in a single cube. The experience from the IBD analysis suggests that the most efficient way to select the positron annihilation cube (AC) is to look for the cube with the most energy in the event. The normalised distribution of the energy found in the AC candidate is shown in Figure 12. The distribution contains two groups of events: with (green) and without (red) pair. The latter shows expectedly two Compton edges issued by the 4.44 MeV gamma from AmBe and the 2.2 MeV gamma from -capture on hydrogen. It also contains proton recoil signals coming from the scattering with AmBe neutrons that are energetic enough, unlike the IBD, to trigger this physics process.
@@
The green signal distribution in Figure 12 shows a sharp peak at 3.4 MeV of energy deposited by the conversion pair. A small shoulder on the right-hand side of this peak is observed and corresponds to additional energy deposits from the annihilation gammas experiencing Compton scattering in the AC. The topological selection of calibration candidates of two back-to-back photons separated by several cubes (as performed with IBD signal candidates discussed in [7]) removes most of this contribution.
@@
This Geant4 simulation is processed with the SoLid reconstruction software Saffron2. Basic physics information about the event, such as the additional cubes in the envelope, the number of reconstructed annihilation gamma candidates, and the cosine of the angle between the annihilation gamma candidates, is built 222The processing is done with the calibration based on the use of the horizontal muons and within the special module called the ReadOut Simulation or ROSim. The goal of this module is to make the Geant4 output files compatible with the Saffron2 input format. ROSim transforms the Geant4 energy deposits into the number of scintillation photons that are read out by the correspondent MPPCs. However, it skips the actual optical simulation and generation of the waveforms to optimise the computational cost of the processing. The detailed tuning of the ROSim is reported in [15].. Both signal and background events are preselected according to the geometric requirements : AC must be a unique cube in the envelope, the number of reconstructed annihilation gamma candidates must be 2, and they must be located in different hemispheres of the detector.
@@
Figure 13 shows the stacked signal (green) and background (red) distributions in the case of Geant4 simulation (left) and ROSim (right). One can notice that the imposed geometrical requirements removed the annihilation gamma-knee energy deposits. The energy resolution of the detector results in a smearing of the measured distribution. The most probable values of both the signal and the background are found in the calibration region of interest.
@@ 5.1.1 Calibration method
Template models based on the Geant4 energy distributions for the signal and background are used to interpret the AmBe calibration data following a method discussed in Section 3.4.2 in [11]. The original model described in the previous section is first smeared event-by-event with a normal distribution. The parameter of the Gaussian represents the energy resolution of the detector. Each resulting distribution is further shifted by an absolute energy scale factor.
The calibration proceeds via a scan of the parameter space for both the and the absolute energy scale factor to find the best pair of parameters. An unbinned event-by-event maximum likelihood fit is performed with the ROOT RooFit toolkit software [16]. For each pair, the transformed templates are converted to probability density functions, while the signal and background yields remain the free parameters of the fit. The negative logarithmic likelihood is used as a measurement of the goodness of fit. The method has been tested with ROSim data samples, separately for signal and background candidates first (see Figure 14) and then by combining them. The best fit is obtained for the pair of parameters (0.932, 14.9%) and is reported in Figure 15.
@@ 5.1.3 Results
The same fit strategy used for the simulation is applied to the calibration data obtained in the 2020 summer calibration campaign. An additional requirement based on the low multiplicity of the cubes (less than 5) is applied (and by construction an event must have at least 3 cubes). Indeed, events with low multiplicity only accommodate the AC and 1 or 2 cubes for each annihilation gamma. This requirement therefore mitigates the contributions of proton recoils in the case of events with multiple neutrons. The best-fit result is shown in Figure 18. The value of the energy scale factor relative to the muon absolute calibration reads .
@@ 6 Crosscheck and validations
Note that the results reported in this section use both relative and absolute calibration performed with the horizontal muons. To begin with, one of the requirements of the calibration was a statistical precision at the level of 1%. To reach it, the ROff data are gathered in 10-day periods. This sample size allows the precision of the method and the actual variations of the detector response to be aligned. Figure 19 shows an example of the statistical uncertainty distribution. The blue value corresponds to an uncertainty associated with one of the fibres in the main cube for each of the detection cells of the SoLid detector, whereas the orange distribution shows the same for one of the fibres impacted by the LL.
@@
Horizontal muons provide information on the time evolution of the detector response. The PVT is exposed to ageing, which decreases the performance of scintillating photon generation. Ageing is not homogeneous; hence, some of the fibres can be more exposed to it. Finally, the WLS fibres also undergo ageing, but the scale of the effect is much less significant, according to the manufacturer’s information. All of these factors imply that the characteristics of light sharing evolve over time. To demonstrate it, we take one of the vertical fibres in the main cube of each detection cell. The fraction of light received by this fibre during the reference period is compared with the fraction received by the same fibre in the following periods. The plot on the left of Figure 20 shows this comparison for the reference period from the very beginning of data-taking in summer 2018 to periods 6, 12, and 18 months later. The mean value of the distribution that shifts towards the negative one is a clear measurement of the ageing of the detector. Simultaneously, the growing width of the distribution shows that ageing impacts the fibres differently; hence the importance of the frequent relative calibration of the detector. The same behaviour is obtained for the four main fibre distributions. The absolute energy scale can be used as an alternative approach to track the evolution of the detector response. The plot on the right of Figure 20 shows the evolution of the absolute energy scale factor. This is yet another demonstration of the ageing of PVT. The average effect is around 4-5% per year, which is slightly higher than expectation [17]. It could be explained by a chiller failure that occurred in May 2019, when the temperature in the detector container increased to 50∘C instead of the usual 12∘C in standard operation. The incident is believed to have caused a degradation in the performance of the PVT. The final cross-check is done with a higher granularity. It explores the evolution of the response on the per-module level. The result is presented in Figure 21. The last detector module concentrates most of the dead fibres; thus, the average absolute energy scale in the module is expected to be lower, while it is determined with slightly worse precision. This effect is mitigated if planes beyond plane 45, where most of the dead fibres reside, are discarded. Without these planes, the performance of module 5 matches the performance of other modules.
@@ 6.1 Cosmogenic background determination
12B is a well-known radioactive source. The half-life of the isotope is short (20.2 ms), and the end-point energy of subsequent decay is 13.4 MeV. Therefore, the energy distribution of these electrons completely covers the IBD energy range of interest. The selection of event candidates is based on the identification of muon tracks that stop inside the detector. Then an electromagnetic cluster, close in space and time to the reconstructed muon, is searched for. To suppress the background, the energy of the cluster is required to be above 3 MeV. Furthermore, events with additional muon candidates within the 200 s time window are removed. It suppresses spallation neutrons that can mimic an electromagnetic cluster. The yield of 12B is estimated from a fit of the time distance between the muon and the electromagnetic cluster. Finally, the energy spectrum reported in Figure 22 is determined after a statistical subtraction of the background using the sPlot [18] technique. The comparison between the measured spectrum and the reference from Geant4 shows an agreement better than 5% throughout the energy range. These results validate both relative and absolute calibration with horizontal muons. This standard candle background could not be observed in the SoLid data prior to the introduction of the muon calibration. The interested reader can find a comprehensive description of this analysis in [19].

### v4

@@ 4 Relative calibration
As discussed in Section 2, a sample of horizontal muons is considered to obtain the system matrix elements that define a cell-by-cell calibration. For each fibre, projection is calculated by multiplying projector by the energy deposit made in the cube . The definition of involves the multiplication of three factors: the MPPC gain (), the absolute energy scale (, which shows the number of photoavalanches generated per unit of energy) and the fraction of light received by the fibre (). The result of the multiplication of by is the total number of scintillation photons (denoted as ) in the given cube, which matches the number of photons in the given plane (since, by construction, horizontal muons hit a unique cube in the plane). The gain for the entire SoLid detector was equalised at the beginning of each data-taking period [12]. As discussed in Section 5, the absolute scale factors are homogenised for all detection cells. Therefore, it is equivalent (up to a scaling factor) to fill the system matrix with the values of or . We arbitrarily chose the latter. In summary, the projection is defined as:
where represents the number of scintillating photons reaching the MPPC . The sum of all projections in the given plane (denoted as ) is the only missing piece to compute . It is expressed as:
Finally, the projector is given by:
@@ 4.1 KL divergence
The left panel of Figure 8 shows the raw distribution of the values for one of the fibres in the adjacent cube (due to LL). A fit to this distribution with a Landau function would allow us to determine the values. However, the presence of outliers compromises the goodness of fit. There are multiple approaches to address outliers, and the Kullback–Leibler (KL) divergence method [13] is one of them. The KL divergence method is a type of statistical distance (or dissimilarity test), which quantifies how much a probability distribution is different from a reference probability distribution . In the case of the SoLid experiment, both distributions must obey the Poisson law; hence, the KL divergence, slightly modified with respect to Ref. [13], is given by:
This class of equations is solved with a Lambert function [14]. This is a multivalued function with different branches. In our implementation, is either 22.5% (main cube with four fibres) or 1.25% (adjacent cubes, 8 fibres). The parameter is the mean of a Gaussian model fit to the initial distribution of . Equation (4.4) is solved for (mean value of the measured distribution) as follows:
As a result, it is not the distribution of values that is fit. Instead, we fit the statistical distance for each set of projector values from the Poisson law. The outliers in the transformed distribution are much more pronounced and can be straightforwardly excluded from the fit. The result of the application of the KL divergence method is shown in Figure 8. As such, the total sum of the light-sharing fractions from the four fibres of the main cube and the eight fibres of the adjacent cubes improves from 96% to 99%.
@@ 4.2 Homogenisation of the response
The second stage of the relative calibration is the homogenisation of the response from individual cubes. The characteristics of the detection units (namely the absolute energy scale ) differ, causing a variation in the response of the detector. The energy loss of the muon serves as a calibration reference. The sum of the projections of the 12 fibres involved within the plane is considered as the representation of the energy deposit made by the muon while crossing it (as in the variable from Equation (4.2)). In addition, all muon candidates are selected from Type 2 muons, for which there is a successfully reconstructed track. The path length of the muon within a certain plane is determined from this track information and is further denoted as . Thus, the energy loss is estimated as:
The energy loss distribution for a given cube is obtained by applying Equation (4.6) to the subset of muons that cross it. A muon is rejected if the cube is not reconstructed with 4 fibres (except for the cases where the fibre is permanently dead). This additional selection compensates for the dying fibre effect presented in Section 3. The result of the selection is shown in Figure 9. A Landau function convoluted with a Gaussian is fit to the resulting energy loss distribution. The model has 3 free parameters: the Gaussian width, the Landau scale factor, and its most probable value (, where shows that the fit result comes from the energy deposit ). The latter is used as the denominator in the homogenisation procedure.
The mean detector response is used as the numerator. It is denoted as and is calculated as the mean value of the distribution with all the values of of the individual detection cubes. Projector values are weighted with respect to the average detector response. As a result of the re-weighting, the values are aligned for all the detection cells. It concludes the relative calibration of the detector and establishes the final values of the elements of the System Matrix. The impact of the homogenisation procedure is, as expected, the most pronounced for planes with permanently dead fibres. As an illustration, Figure 10 displays the response of the 4 fibres of each main cube in plane 48. The dark blue area (half of the plane) corresponds to the dead fibres. Even under such conditions, the relative calibration works reasonably well and allows the events appearing in that region of the detector to be adequately reconstructed.
@@ 5 Absolute calibration
As a result of the relative calibration, the absolute energy scale is homogenised for all the detection units but remains unknown. The absolute energy scale is a factor (denoted as in Equation (4.1)) that converts the energy measurement from the units provided by the electronics (ADC) to physical units (MeV). The determination of the (now) universal scale factor completes the electromagnetic calibration of the detector.
5.1 Absolute calibration with the horizontal muons
One possibility of determining the absolute energy scale consists of using the energy-loss distribution from the cosmic horizontal muon. The energy loss values for all the detection cells are gathered in a single distribution to which the Landau function convoluted with the Gaussian is fit. The most probable value of Landau gives the reference value in ADC units 111Technically in pixel avalanches, or PA, but since the gain is equalised for all the fibres, the conversion factor from PA to ADC is known.. The value in MeV is obtained from a dedicated Geant4 simulation. It includes the full-scale detector geometry that is combined with the CRY engine [15] to simulate muons. All energy loss values obtained from the simulation are also combined in a single distribution. The ratio between the most probable values extracted from two fits defines the absolute energy scale.
Figure 11 shows the distributions of the energy loss values obtained with muons in three different energy regimes (0.5, 5, and 50 GeV). Variations in the average response are at the level of 10%. This provides the fundamental limit to the precision of this method. Any improvement would require a more accurate knowledge of the energy distribution of horizontal muons.
@@ 5.2 Absolute calibration with the americium-beryllium source
A novel method (to our knowledge) has been introduced by making use of an americium-beryllium (AmBe) radioactive calibration source that was used in SoLid to calibrate the neutron signal. The idea consists of measuring the response of the detector to the 4.44 MeV gamma emitted by excited 12C, with a branching fraction of 60%. These gammas can experience an electron-positron () pair conversion that provides a well-defined monoenergetic calibration source of 3.4 MeV. Although Compton scattering is the dominant process, the conversion appears in 3.5% of the cases according to a generic Geant4 simulation (with the G4EmStandardPhysics physics list). The presence of the positron makes the calibration signature identical to that of the IBD signal candidates. Therefore, all tools developed for the topological identification of IBD events are highly relevant and can be used to select .
The chain that produces 4.4 MeV gamma rays is as follows:
@@
The initial information on the electromagnetic part of the signal provided by AmBe is obtained from a dedicated Geant4 simulation. It consists of a simulation of the full-scale SoLid detector geometry, where the radioactive source is placed in one of the positions available between the modules. To optimise the computational cost of the process, only and are simulated directly, instead of the entire chain shown in Equation (5.1). The simulation indicates that the pair deposits its energy predominantly in a single cube. The experience from the IBD analysis suggests that the most efficient way to select the positron annihilation cube (AC) is to look for the cube with the most energy in the event. The normalised distribution of the energy found in the AC candidate is shown in Figure 12. The distribution contains two groups of events: with (dashed blue) and without (solid orange) pair. The latter shows expectedly two Compton edges issued by the 4.44 MeV gamma from AmBe and the 2.2 MeV gamma from -capture on hydrogen. It also contains proton recoil signals coming from the scattering with AmBe neutrons that are energetic enough, unlike the IBD, to trigger this physics process.
@@
The dashed blue signal distribution in Figure 12 shows a sharp peak at 3.4 MeV of energy deposited by the conversion pair. A small shoulder on the right-hand side of this peak is observed and corresponds to additional energy deposits from the annihilation gammas experiencing Compton scattering in the AC. The topological selection of calibration candidates of two back-to-back photons separated by several cubes (as performed with IBD signal candidates discussed in Ref. [8]) removes most of this contribution.
@@
This Geant4 simulation is processed with the SoLid reconstruction software Saffron2. Basic physics information about the event, such as the additional cubes in the envelope, the number of reconstructed annihilation gamma candidates, and the cosine of the angle between the annihilation gamma candidates, is built 222The processing is done with the calibration based on the use of the horizontal muons and within the special module called the ReadOut Simulation or ROSim. The goal of this module is to make the Geant4 output files compatible with the Saffron2 input format. ROSim transforms the Geant4 energy deposits into the number of scintillation photons that are read out by the correspondent MPPCs. However, it skips the actual optical simulation and generation of the waveforms to optimise the computational cost of the processing. The detailed tuning of the ROSim is reported in Ref. [16].. Both signal and background events are preselected according to the geometric requirements : AC must be a unique cube in the envelope, the number of reconstructed annihilation gamma candidates must be 2, and they must be located in different hemispheres of the detector.
@@
Figure 13 shows the stacked signal (empty blue) and background (orange diagonal cross-hatch) distributions in the case of Geant4 simulation (left) and ROSim (right). One can notice that the imposed geometrical requirements removed the annihilation gamma-knee energy deposits. The energy resolution of the detector results in a smearing of the measured distribution. The most probable values of both the signal and the background are found in the calibration region of interest.
@@ 5.2.1 Calibration method
Template models based on the Geant4 energy distributions for the signal and background are used to interpret the AmBe calibration data following a method discussed in Section 3.4.2 in Ref. [12]. The original model described in the previous section is first smeared event-by-event with a normal distribution. Namely, each cube energy is randomly varied following a standard distribution with = as indicated by 22Na calibration results [12]. Each resulting distribution is further scaled by an absolute energy scale factor.
The calibration proceeds via a scan of the parameter space for both the and the absolute energy scale factor to find the best pair of parameters. An unbinned event-by-event Maximum Likelihood (ML) fit is performed with the ROOT RooFit toolkit software [17]. For each pair, the transformed templates are converted to probability density functions, while the signal and background yields remain the free parameters of the fit. The negative logarithmic likelihood is used as a measurement of the goodness of fit. The method has been tested with ROSim data samples, separately for signal and background candidates first (see Figure 14) and then by combining them. The best fit is obtained for the pair of parameters (0.932, 14.9%) and is reported in Figure 15.
@@ 5.2.3 Results
The same fit strategy used for the simulation is applied to the calibration data obtained in the 2020 summer calibration campaign. This campaign (as well as any other that involves the radioactive sources) in the SoLid experiment is performed with the CROSS (CalibRation On Site Solid) robotic hand. It is capable of inserting the source between the module and placing it in one of the nine predefined positions. The calibration data used in this paper was obtained with the source located between the last two modules at four different locations (central, lower left, lower right, and upper right). The total data-taking time was 77 hours. An additional requirement based on the low multiplicity of the cubes (less than 5) is applied (and by construction, an event must have at least 3 cubes). Indeed, events with low multiplicity only accommodate the AC and 1 or 2 cubes for each annihilation gamma. This requirement, therefore, mitigates the contributions of proton recoils in the case of events with multiple neutrons. The best-fit result is shown in Figure 18. The value of the energy scale factor relative to the muon absolute calibration reads .
@@ 6 Crosscheck and validations
Note that the results reported in this section use both relative and absolute calibration performed with the horizontal muons. To begin with, one of the calibration requirements was a statistical precision at the level of 1%. To achieve it, the ROff data is gathered in 10-day periods. This sample size allows the precision of the method and the actual variations of the detector response to be aligned. Figure 19 shows an example of the statistical uncertainty distribution. The blue value corresponds to an uncertainty associated with one of the fibres in the main cube for each of the detection cells of the SoLid detector, whereas the orange distribution shows the same for one of the fibres impacted by the LL.
Horizontal muons provide information on the time evolution of the detector response. The PVT is exposed to ageing, which decreases the performance of scintillating photon generation. Ageing is not homogeneous; hence, some of the fibres can be more exposed to it. Finally, the WLS fibres also undergo ageing, but the scale of the effect is much less significant, according to the manufacturer’s information. All of these factors imply that the characteristics of light sharing evolve over time.
To demonstrate it, we take one of the vertical fibres in the main cube of each detection cell. The fraction of light ( value) received by this fibre during the reference period is compared with the fraction received by the same fibre in the following periods. The - axis in Figure 20-left shows the evolution of the spread of the value by computing the fraction . Where corresponds to the period from the very beginning of data-taking in summer 2018, and are the periods 6, 12, and 18 months later, respectively. The mean value of the distribution that shifts towards the negative one is a clear measurement of the ageing of the detector. Simultaneously, the growing width of the distribution shows that ageing impacts the fibres differently; hence the importance of the frequent relative calibration of the detector. The same behaviour is obtained for the four main fibre distributions.
@@
The absolute energy scale can be used as an alternative approach to track the evolution of the detector response. The plot on the right of Figure 20 shows the evolution of the absolute energy scale factor. This is yet another demonstration of the ageing of PVT. The average effect is around 4 - 5% per year, which is slightly higher than expectation [18]. It could be explained by a chiller failure that occurred in May 2019, when the temperature in the detector container increased to 50∘C instead of the usual 12∘C in standard operation. The incident is believed to have caused a degradation in the performance of the PVT. The final cross-check is done with a higher granularity. It explores the evolution of the response on the per-module level. The result is presented in Figure 21. The last detector module concentrates most of the dead fibres; thus, the average absolute energy scale in the module is expected to be lower, while it is determined with slightly worse precision. This effect is mitigated if planes beyond plane 45, where most of the dead fibres reside, are discarded. Without these planes, the performance of module 5 matches the performance of other modules.
@@ 6.1 Cosmogenic background determination
12B is a well-known radioactive source. The half-life of the isotope is short ( 20.2 ms), and the end-point energy of subsequent decay is 13.4 MeV. Therefore, the energy distribution of these electrons completely covers the IBD energy range of interest. The selection of event candidates is based on the identification of muon tracks that stop inside the detector. Then an electromagnetic cluster, close in space and time to the reconstructed muon, is searched. To suppress the background, the energy of the cluster is required to be above 3 MeV. Furthermore, events with additional muon candidates within the 200 s time window are removed. It suppresses spallation neutrons that can mimic an electromagnetic cluster. The yield of 12B is estimated from a fit of the time distance between the muon and the electromagnetic cluster. Finally, the energy spectrum reported in Figure 22 is determined after a statistical subtraction of the background using the sPlot [19] technique. The comparison between the measured spectrum and the reference from Geant4 shows an agreement better than 5% throughout the energy range. These results validate both relative and absolute calibration with horizontal muons. This standard candle background could not be observed in the SoLid data prior to the introduction of the muon calibration. The interested reader can find a comprehensive description of this analysis in Ref. [20].